Calculate the decrease in available energy when 25 kg of water at 95oC mix with 35 kg of water at 35oC, the pressure being taken as constant and the temperature of the surroundings being 15oC. (Cp of water = 4.2 kJ/kg K).b.Find entropy, enthalpy and specific volume of steam at 173oC.


1
Expert's answer
2021-07-29T17:21:02-0400

Entropy


"s_1=mc_p(T_2-T_1)=25\times4.2\times(368-308)=6300KJ/K"

"s_2=mc_p(T_2-T_1)=35\times4.2\times(368-308)=8820KJ/K"

"\u2206s=s_2-s_1=8820-6300=2520kJ\/kg"

Enthalpy

Phase change


"\u2206H=H_2-H_1=4.2(368-308)=452kJ\/kg"

Specific volume

"V=\frac{v}{m}\\V=\frac{173}{25+35}=\frac{173}{60}=2.88m^3/kg"